OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given point and having the given slope.

OpenStudy (jack1):

straight line : y = mx + c quadratic line: y = ax^2 + bc + c circular line: x^2 + y^2 = 1

OpenStudy (anonymous):

E(3, 5), m = undefined can you help me with this? this is from the first question

OpenStudy (jack1):

you need at least 2 points to work out the equation of a straight line, sorry dude

OpenStudy (amistre64):

standard form is not the same as slope intercept form

OpenStudy (jack1):

..."standard ...form..."...u say?

OpenStudy (amistre64):

and you only need a slope and a point to define a line. 2 points helps to define a slope is all

OpenStudy (amistre64):

standard from is: Ax + By = C

OpenStudy (jack1):

but we dont have slope...?

OpenStudy (amistre64):

m = undefined is the slope

OpenStudy (amistre64):

what type of line has an undefined slope?

OpenStudy (anonymous):

jack, thats all the question gives me tho

OpenStudy (jack1):



OpenStudy (amistre64):

solving the stadard form for y, we get the slope intercept form \[ax+by=c\]\[y=-\frac abx+\frac cb\] given that m = -a/b is undefined would suggest that b=0 Ax + 0y = C Ax = C is the line of an undefined slope

OpenStudy (amistre64):

so yes

OpenStudy (amistre64):

using the point (3,5) Ax = C A(3) = C , when C/A = 3 in simplest form A would equal 1 giving us: C = 3 therefore the equation of line of an undefined slope at (3,5) is: x = 3